write down kapler third law for timeperiod
1
Expert's answer
2020-09-17T14:29:18-0400

Solution

According to Kepler's third law states that square of

Time period of an orbiting planet around sun is directly proportional to cube of semi major axis.


"T^2=\frac{4\pi^2a^3}{GM}"

Where

a is semi major axis.

G is gravitational constant.

M is mass of sun.